# A Survey on Bundle Recommendation: Methods, Applications, and Challenges

### v1

@@ 4.4. Summary of Existing Generative BR Research
(Chen et al., 2019a; Yang et al., 2024b; Interdonato et al., 2013; Beladev et al., 2016)
@@ 5.1. Datasets in Various Domains
We summarize 30 datasets used for bundle recommendation research across 11 application domains. These datasets are listed in Table 8 and categorized by domain including music playlist, book list, movie list, outfit, meal, drug package, service package and so on. Table 8 illustrates the domains, sources, the type of data each dataset contains, and how the bundle in each dataset is presented.
@@
Domain Name of Dataset Source Data Description of Bundle Available or Not References
Playlist NetEase Netease Cloud Music
users,songs,playlist
@@
A playlist contains user-defined collections of songs
on a specific theme.
@@
users,items,lists
user-list,density
@@
AOTM Art of the Mix
users,items,lists
list-item
avg. list interactions
@@
users,books,lists
user-list,user-book,list-book
avg. list interactions
avg. book interactions
@@
users,books,lists
user-book,user-list
@@
Movie MovieLens published by GroupLens research group
users,movies
user-movie
@@
users,items,lists
@@
users,lists,items
user-list,user-item
@@
Outfit iFashion the online application of Taobao
users,items,outfits
@@
Clothing Amazon
users,items,bundles
@@
categories,density
A bundle consists of several clothes. Y (Sun et al., 2022b; He et al., 2022)
@@
from an online bundle list
recommendation service of
a commercial e-business, Taobao
users,items,bundles
categories,records
@@
Polyvore Polyvore, a well-known fashion website
users,items,outfits
@@
multimodal information
@@
Game List Steam Steam video game distribution platform
users,games,bundles
user-game,user-bundle
@@
A bundle is a group of different games sold together on website. Y
@@
the mobile game Love is Justice
developed by Netease Games
users,items,bundles
@@
A bundle is made up of props(virtual items) in the game. - (Deng et al., 2020)
Games
NetEase Games’ massively
multiplayer online role-playing game
users,items,bundles
avg. bundle interactions
avg. item interactions
@@
A bundle consists of the virtual items sold in the game. - (Li et al., 2023c)
@@
Allrecipes, one of the most popular
food sharing sites in the world
users,meals,courses
user-meal,user-course,meal-course
A meal contains multiple courses from specific categories. Y (Li et al., 2022a, 2024a, 2023a)
Travel Package Tuniu
@@
users,POIs,trajectories
POI visits
@@
users,POIs,trajectories
POI visits
@@
POI check-ins,pacakge check-ins
avg. POI check-in per user
@@
Drug Package APH
from the electronic medical
record database of a
first-rate hospital in China
patients,medical records
disease document,drugs
drug packages
interactions between drugs
@@
patients,medical records
disease document,medicines
medicine packages
@@
Service
Package
ProgrammableWeb
@@
services,mashups
functional description
@@
QA List Zhihu a question and answer community
users,items,lists
user-list
density
@@
Product Bundles cosmestics from one online merchant in Taobao
users
skin care products
transactions
A product bundle consists of two or more items
transacted by users.
@@
perfume from one online merchant in Taobao
users
perfumes
transactions
A product bundle consists of two or more items
transacted by users.
@@
TaFeng a Chinese grocery store dataset
@@
Dunnhumby a household-level transaction dataset
@@
(Li et al., 2023d; Deng et al., 2023a)
Instacart
an online grocery shopping dataset
@@
Music. A playlist usually consists of songs on specific theme. The typical one is NetEase constructed by (Cao et al., 2017). They have collected data from Netease Cloud Music, which enables users to construct a list of songs with a specific theme. Another commonly utilized dataset is Spotify dataset (He et al., 2019b, 2020b). The playlists are created by users with some popular topics. Additionally, AOTM dataset, proposed by (McFee and Lanckriet, 2012), was created by collecting playlists from Art of the Mix, a website where users can upload their song playlists.
@@
Book. A booklist contains multiple books with different genres. One prominent dataset in this domain is named Youshu, provided by (Chen et al., 2019b), which is constructed by crawling data from Youshu, a Chinese book review site. Similarly, Goodreads dataset which serves as a comprehensive repository of book-related information has been leveraged in studies (He et al., 2019b, 2020b; Liu et al., 2014), comprising user-generated reviews, ratings, and social interactions.
@@
Movie. A movie list is a set of movies. One notable dataset in this domain is the MovieLens [^1] https://movielens.org/, collected rating data sets from the MovieLens web site by GroupLens research group. Since there are only interactions with individual movies in the dataset, bundle recommendation systems usually package movies into lists and then recommend lists to users (Xie et al., 2010; He et al., 2022; Kabutoya et al., 2013). Another dataset is Douban, crawled from an online social platform Douban by (Yang et al., 2020), which enables users to select their interested movies or user generated movie watching lists. MX Player, one of the largest streaming platforms in India, organizes and exposes video via lists. There are many studies (Li et al., 2021a, b) that constructed their own datasets from the MX Player platform to conduct experiments.
@@
Outfit. An outfit is a set of well-matched fashion items, a form of bundle. The iFashion dataset constructed by (Chen et al., 2019a) serves as a common dataset used in bundle recommendation (Ma et al., 2022; Ren et al., 2023; Du et al., 2023; Ma et al., 2024a; Jeon et al., 2024; Liu et al., 2023b). Another one, Clothing is a public bundle dataset constructed from the Amazon e-commerce dataset (Sun et al., 2022b), where bundles are a combination of several clothes based on user intents. Similarly, Polyvore dataset is collected by (Han et al., 2017) from Ployvore website containing outfits and items (each item contains a pair - product image and a corresponding text description). Taobao dataset, utilized in (Bai et al., 2019), is collected from an online bundle list recommendation service of a commercial e-business, Taobao. The bundles in the Taobao dataset are predefined manually by the sellers.
@@
Meal. A meal contains multiple courses from specific categories. The MealRec dataset, proposed by (Li et al., 2022a), is a publicly accessible dataset tailored for meal recommendation research. It is compiled from the user review records of Allrecipe.com, including users, recipes and meals. Due to the lack of publicly available meal recommendation datasets including meal-course affiliation, MealRec+, a new meal recommendation benchmark dataset, is introduced by (Li et al., 2024a). It fills a data gap by including meal-course affiliation and healthiness information.
@@
Game. A game list is a group of different games sold together on website. The Steam dataset is collected from the Steam video game distribution platform by (Pathak et al., 2017), including gamers, game bundles, purchase records for both individual games and bundles. Moreover, the Justice, an industrial dataset, is collected from the mobile game Love is Justice developed by Netease Games, where bundles are made up of virtual items in the game. One dataset named Games(Li et al., 2023c) is also collected from NetEase Games, a massively multiplayer online role-playing game, whose bundles consist of the virtual items sold in the game.
@@
Travel. A travel package is an ordered set of multiple POIs with different types, where the order corresponds to the visiting sequence. The Flickr and Foursquare datasets are widely used in (Chen et al., 2016; Wang et al., 2023a; Gao et al., 2022; He et al., 2019a). Flickr dataset originates from the popular image-sharing platform Flickr. Foursquare dataset comes from the location-based social networking service Foursquare. TripAdvisor dataset, widely used in (Xie et al., 2010; Benouaret and Lenne, 2016, 2017), is derived from the well-known travel website where users can share and explore travel information. They both contain user interactions with POIs such as reviews, ratings, and check-ins, along with location data. Furthermore, a real check-in dataset obtained from Gowalla (Fang et al., 2014) contains users, POIs, and check-ins. Another real-world travel dataset is provided by Tuniu, one of the largest tourism e-commerce platforms in China. It is comprised of page-to-page clickstream data recorded in the server logs, which has been adopted in the studies (Zhu et al., 2023a, 2021).
Drug. A drug package usually consists of several drugs with drug interactions, such as synergism, antagonism or no interaction. There are two medical datasets, APH (Zheng et al., 2023) and MIMIC-III (Johnson et al., 2016). The data of patients with hospital visit are regarded as medical record. And each record contains the disease description documents of the patients. In general, the dataset includes medical records of patients, involving disease documents, drugs, and drug packages.
@@
Service. A service package, also called API package, is a bundle of compatible services, which can satisfy the functional requirements of the mashup as completely as possible. The real-world ProgrammableWeb dataset, which is widely used in existing service bundle recommendation studies (Gu et al., 2021; Liu et al., 2021, 2023c; Liu and Cao, 2020; Qi et al., 2021; Gu et al., 2016; Cao et al., 2016), is collected from ProgrammableWeb.com, a largest mashup information sharing community. The dataset generally contains services (called APIs), mashups and some side information, such as descriptions which describe mashup and service functionalities, tags, etc.
@@
Others. Additionally, several other datasets contribute to the field of bundle recommendation research. One such dataset is Zhihu, constructed by (He et al., 2019b), which gathers data from Zhihu, a question and answer community where users can follow lists of answers that have been curated by other users. Furthermore, TaFeng, Dunnhumby, and Instacart datasets are widely used for next basket recommendation studies (Yu et al., 2023; Sun et al., 2023b; Su et al., 2024; Deng et al., 2023b).
@@ 5.2. Evaluation Metrics
As shown in Table 9, it illustrates some of the key metrics commonly used in the evaluation of bundle recommendation systems (Chang et al., 2020; Chen et al., 2019b; Zhao et al., 2022; Ma et al., 2022, 2024a; Li et al., 2024b; Yang et al., 2020), along with their formulas and brief descriptions. We summarize 4 metrics for accuracy, 3 metrics for ranking quality, 4 metrics for diversity, 2 metrics for ranking exposure, and other metrics like AUC. These metrics provide a comprehensive understanding of how well the system retrieves and ranks relevant bundles. Given that $\mathbb{I}(\cdot)$ is the indicator function, which is 1 if the condition inside is true, and 0 otherwise; threshold is a predefined value used to determine if a prediction is considered positive; $\text{rank}(b_{i})$ is the predicted rank of the bundle $b_{i}$ for user $u$; $\text{sim}(b_{i},b_{j})$ is a similarity measure between two bundles $b_{i}$ and $b_{j}$; $\hat{y}_{ub}$ is the predicted score, and $y_{ub}$ is the actual score; $\mathbf{R}(k)$ is a list of k recommended bundles; $app(i,\mathbf{R}(k))$ indicates whether item $i$ appears in $\mathbf{R}(k)$; $p(i,\mathbf{R}(k))$ indicates the sum of item $i$’s frequency in user $u$’s recommended bundles; $utility(x\mid r)$ denotes the utility function, which is usually a monotonically decreasing function with respect to the rank of a bundle. One common choice is to use an exponent $w>=0$, and $utility(x\mid r)=[\operatorname{rank}(x\mid r)]^{-w}$.
@@
Entropy $-\sum_{i\in\mathcal{I}}p(i,\mathbf{R}(k))\log p(i,\mathbf{R}(k))$

### v2

@@ 4.4. Summary of Existing Generative BR Research
(Chen et al., 2019b; Yang et al., 2024b; Interdonato et al., 2013; Beladev et al., 2016)
@@ 5.1. Datasets in Various Domains
We summarize 30 datasets used for bundle recommendation research across 11 application domains. These datasets are listed in Table 9 and categorized by domain including music playlist, book list, movie list, outfit, meal, drug package, service package and so on. Table 9 illustrates the domains, sources, the type of data each dataset contains, and how the bundle in each dataset is presented. Moreover, we also present whether content information, such as text, images, or acoustics, is available in datasets, which provides researchers with valuable entry points for leveraging LLMs to further explore bundle recommendation.
@@
Domain Name of Dataset Source Data Content Information Description of Bundle Available or Not References
NetEase Netease Cloud Music
users,songs,playlist,
@@
A playlist contains user-defined
collections of songs on a specific theme.
@@
users,items,lists,
user-list,list-item,density
music titles, playlist title,
album name and artist names
Spotify URIs and the playlist’s
number of followers
@@
Playlist AOTM Art of the Mix
users,items,lists,
list-item,
avg. list interactions,
@@
users,books,lists,
user-list,user-book,list-book,
avg. list interactions,
avg. book interactions,
@@
users,books,lists,
user-book,user-list,
@@
book title, language,
description, genre,
authors, etc.
@@
MovieLens
published by GroupLens
research group
users,movies,
user-movie ratings
user gender, age, occupation,
movie titles, genres
@@
users,items,lists,
@@
users,lists,items,
user-list,user-item,
@@
iFashion
originating from
the POG dataset
users,items,outfits,
@@
Clothing originating from Amazon
users,items,bundles,
@@
bundle intents, category, title A bundle consists of several clothes. Y (Sun et al., 2022b; He et al., 2022)
@@
from an online bundle
list recommendation
service of a commercial
e-business, Taobao
users,items,bundles,
categories,records,
@@
Outfit Polyvore
Polyvore, a well-known
fashion website
users,items,outfits,
@@
fashion item descriptions, images,
number of likes of the outfit,
hash tags of the outfits, etc.
@@
Steam
Steam video game
distribution platform
users,games,bundles,
user-game,user-bundle,
@@
bundle name, bundle price,
bundle url, genre, item name,
item url, pricing information
A bundle is a group of different
games sold together on website.
Y
@@
the mobile game
Love is Justice
developed by
Netease Games
users,items,bundles,
@@
A bundle is made up of props
(virtual items) in the game.
- (Deng et al., 2020)
Game List Games
NetEase Games’
massively multiplayer
online role-playing game
users,items,bundles,
avg. bundle interactions,
avg. item interactions,
@@
A bundle consists of the virtual
items sold in the game.
- (Li et al., 2023d)
@@
Allrecipes, one popular
food sharing site in the world
users,recipes,meals,
user-meal,user-recipe,meal-recipe
recipe name, image, category,
rate, ingredients, nutrition, tags
A meal is a bundle composed of
three recipes from different categories.
Y (Li et al., 2022b, 2024a, 2023b)
Tuniu
@@
POI thematic category,
the average rating score,
the number of ratings, etc.
@@
users,POIs,trajectories,
user-POI visits
POI photo, longitude, latitude, theme,
dateTaken, frequency, travel sequence
@@
users,POIs,trajectories,
user-POI visits
time stamp, GPS coordinates,
venue-categories, tag
@@
user-POI check-ins,
user-pacakge check-ins,
package-POI,
avg. POI check-in per user,
@@
POI category, stay time, cost,
package total cost, total stay time,
distances between POIs,
user features
@@
EMR
from the electronic
medical record database
of a first-rate hospital in China
patients, visit records, demographics,
labs, prescriptions, drug packages,
explicit drug–drug interactions
admission notes,
disease description documents
@@
patients, admissions, ICU stays,
vitals, labs, medications, procedures,
diagnostic codes, survival outcomes
clinical notes, discharge summaries,
imaging reports,
physiological waveforms
@@
Service Package ProgrammableWeb
@@
mashups,services,
mashup-service,
avg. service per mashup
functional description, name,
url, category, format, provider,
developer, etc.
@@
QA List Zhihu
a question and
answer community
users,answers,lists,
user-list,list-answer
answer content, question semantics,
list semantics, topics
@@
cosmestics
from one online
merchant in Taobao
customers,skin care products,
user-item transaction records,
user-bundle transaction records
product name, brand, price,
functionality, origin, fragrance,
transaction price variance,
price distribution
A product bundle consists of two or more items transacted
by users.
@@
perfume
from one online
merchant in Taobao
customers,perfumes,
user-item transaction records,
user-bundle transaction records
product name, brand, price,
functionality, origin, fragrance,
transaction price variance,
price distribution
A product bundle consists of two or more items transacted
by users.
@@
TaFeng
a Chinese grocery
store dataset
@@
transaction date and time,
age group, PIN code, product subclass,
amount, asset, sales price
@@
Dunnhumby
a household-level
transaction dataset
@@
description for the campaign,
marital status, income description,
product department, brand,
customer age
@@
(Li et al., 2023c; Deng et al., 2023a)
Co-consumed Products Instacart
an online grocery
shopping dataset
@@
aisle category,
department category,
item name
@@
Music. A playlist usually consists of songs on specific theme. The typical one is NetEase constructed by (Cao et al., 2017). They have collected data from Netease Cloud Music, which enables users to construct a list of songs with a specific theme. Another commonly utilized dataset is Spotify dataset (He et al., 2019b, 2020b). Although the Spotify adopted in these studies are mainly constructed with interaction data, the original Spotify dataset actually contains richer content information, such as playlist titles, artist names, URLs, etc. Additionally, AOTM dataset, proposed by (McFee and Lanckriet, 2012), was created by collecting playlists from Art of the Mix, a website where users can upload their song playlists.
@@
Book. A booklist contains multiple books with different genres. One prominent dataset in this domain is named Youshu, provided by (Chen et al., 2019a), which is constructed by crawling data from Youshu, a Chinese book review site. Similarly, Goodreads dataset which serves as a comprehensive repository of book-related information has been leveraged in studies (He et al., 2019b, 2020b; Liu et al., 2014), comprising book title, language, description, genre, user-generated reviews, ratings, etc.
@@
Movie. A movie list is a set of movies. One notable dataset in this domain is the MovieLens, collected rating data sets from the MovieLens web site by GroupLens research group. Since there are only interactions with individual movies in the dataset, bundle recommendation studies usually package movies into lists and then recommend lists to users (Xie et al., 2010; He et al., 2022; Kabutoya et al., 2013). Another dataset is Douban, crawled from an online social platform Douban by (Yang et al., 2020), which enables users to select interested movies or user generated movie watching lists. MX Player, one of the largest streaming platforms in India, organizes and exposes video via lists that include names and contents. Several studies (Li et al., 2021a, b) have constructed their own datasets from MX Player to conduct experiments.
@@
Outfit. An outfit is a set of well-matched fashion items, a form of bundle. The widely used iFashion dataset (Chen et al., 2019b) provides category, image, and title information, and has served as a benchmark in numerous bundle recommendation studies (Ma et al., 2022; Ren et al., 2023; Du et al., 2023; Ma et al., 2024a; Jeon et al., 2024; Liu et al., 2023b). The Clothing dataset (Sun et al., 2022b), derived from Amazon, is a public bundle dataset in which each bundle consists of several clothes annotated with intents, categories, and titles. The Polyvore dataset (Han et al., 2017), collected from the Polyvore website, contains outfits and items with product images, tags, and textual descriptions. The Taobao dataset (Bai et al., 2019), collected from an online bundle list recommendation service of the e-commerce platform Taobao, contains bundles manually predefined by the sellers.
@@
Meal. A meal contains multiple courses from specific categories. The MealRec dataset, proposed by (Li et al., 2022b), is a publicly accessible dataset tailored for meal recommendation research. It is constructed from user review records of Allrecipe.com, including users, recipes and meals. Each recipe is enriched with detailed attributes such as name, rating, image URL, category, tag, and other metadata. Due to the lack of publicly available meal recommendation datasets including meal-course affiliation, MealRec+, a new meal recommendation benchmark dataset, is introduced by (Li et al., 2024a). It fills a data gap by including meal-course affiliation and healthiness information.
@@
Game. A game list is a group of different games sold together on website. The Steam dataset, collected from the Steam video game distribution platform by (Pathak et al., 2017), contains gamers, game bundles, and purchase records for both individual games and bundles. In addition, the Steam dataset provides rich content information such as bundle names, prices, URLs, genres, game titles, and pricing details. The Justice, an industrial dataset, is collected from the mobile game Love is Justice developed by Netease Games, where bundles consist of virtual items in the game. One dataset named Games (Li et al., 2023d) is also collected from NetEase Games, a massively multiplayer online role-playing game, whose bundles consist of the virtual items sold in the game.
@@
Travel. A travel package is an ordered set of multiple POIs, where the order corresponds to the visiting sequence. The Flickr and Foursquare datasets are widely used in (Chen et al., 2016; Wang et al., 2023b; Gao et al., 2022; He et al., 2019a). Both provide user–POI interactions along with rich content information including geographical coordinates, timestamps, and categorical attributes. Their differences lie mainly in data sources and emphasis: Flickr collects POIs offering additional features like photos, visit frequencies, and travel sequences, whereas Foursquare records explicit user check-ins enriched with venue tags. The TripAdvisor dataset (Xie et al., 2010; Benouaret and Lenne, 2016, 2017), derived from the well-known travel website where users can share and explore travel information, provides user reviews, ratings, and check-ins, with each POI further described by thematic categories, average ratings, and rating counts. The Gowalla dataset (Fang et al., 2014) contains users, POIs, and check-ins, along with content information including attraction categories, attraction cost, package compositions, and distances between attractions. Another real-world travel dataset is provided by Tuniu, one of the largest tourism e-commerce platforms in China. It comprises page-to-page clickstream data recorded in server logs and has been adopted in studies (Zhu et al., 2023a, 2021).
Drug. A drug package refers to a set of drugs prescribed together during a patient’s hospitalization, where drugs may interact through synergism, antagonism, or not at all. Two widely used datasets are MIMIC-III (Johnson et al., 2016) and EMR (Zheng et al., 2023). MIMIC-III provides rich metadata, including patient profiles, admissions, ICU stays, medications, laboratory results, procedure and diagnostic codes, and survival outcomes, as well as content such as clinical notes and imaging reports. In contrast, EMR, constructed from the electronic medical record database of a top hospital in China, emphasizes drug packages with explicit drug–drug interaction annotations, alongside structured records of demographics, lab tests, and prescriptions, as well as admission notes as textual content.
@@
Service. A service package, or API package, is a bundle of compatible services that satisfies the functional requirements of a mashup as fully as possible. The widely used ProgrammableWeb dataset (Gu et al., 2021; Liu et al., 2021, 2023a; Liu and Cao, 2020; Qi et al., 2021; Gu et al., 2016; Cao et al., 2016) is collected from ProgrammableWeb.com, the largest mashup information sharing community. The dataset contains services, mashups, mashup-service relations, and content such as textual descriptions, tags, and categories.
@@
Others. Several other datasets contribute to the field of bundle recommendation research. One such dataset is Zhihu, constructed by (He et al., 2019b), which gathers data from Zhihu, a question and answer community where users can follow lists of answers that have been curated by other users. Furthermore, TaFeng, Dunnhumby, and Instacart datasets are widely used for next basket recommendation studies (Yu et al., 2023; Sun et al., 2023b; Su et al., 2024; Deng et al., 2023b). More details can be found in Table 9.
@@ 5.2. Evaluation Metrics
As shown in Table 10, it illustrates some of the key metrics commonly used in the evaluation of bundle recommendation systems (Chang et al., 2020; Chen et al., 2019a; Zhao et al., 2022; Ma et al., 2022, 2024a; Li et al., 2024b; Yang et al., 2020), along with their formulas and brief descriptions. We summarize 4 metrics for accuracy, 3 metrics for ranking quality, 4 metrics for diversity, 2 metrics for ranking exposure, and other metrics like AUC. These metrics provide a comprehensive understanding of how well the system retrieves and ranks relevant bundles. Given that $\mathbb{I}(\cdot)$ is the indicator function, which is 1 if the condition inside is true, and 0 otherwise; threshold is a predefined value used to determine if a prediction is considered positive; $\text{rank}(b_{i})$ is the predicted rank of the bundle $b_{i}$ for user $u$; $\text{sim}(b_{i},b_{j})$ is a similarity measure between two bundles $b_{i}$ and $b_{j}$; $\hat{y}_{ub}$ is the predicted score, and $y_{ub}$ is the actual score; $\mathbf{R}(k)$ is a list of k recommended bundles; $app(i,\mathbf{R}(k))$ indicates whether item $i$ appears in $\mathbf{R}(k)$; $p(i,\mathbf{R}(k))$ indicates the sum of item $i$’s frequency in user $u$’s recommended bundles; $utility(x\mid r)$ denotes the utility function, which is usually a monotonically decreasing function with respect to the rank of a bundle. One common choice is to use an exponent $w>=0$, and $utility(x\mid r)=[\operatorname{rank}(x\mid r)]^{-w}$.
@@
Entropy $\sum_{i\in\mathcal{I}}p(i,\mathbf{R}(k))\log p(i,\mathbf{R}(k))$